OpenStudy (anonymous):

One of your friends sends you an email asking you to explain how all of the following expressions have the same answer. the cube root of x cubed :x to the one–third power • :x to the one–third power • x to the one–third power 1 over x to the –1 power the eleventh root of the quantity of x to the fifth times x to the fourth times x squared Compose an email back assisting your friend and highlight the names of the properties of exponents when you use them. Can u please help me out? @jim_thompson5910

OpenStudy (anonymous):

\[\sqrt[3]{x^3}=x\] because, well because it is obvious that is what the cubed root means

OpenStudy (anonymous):

I know they all equal x I've already solved for all of them but Idk how to highlight the names of the properties of exponents.

OpenStudy (anonymous):

idk what that's supposed to mean.

OpenStudy (anonymous):

\[\large x^{\frac{1}{3}}\times x^{\frac{1}{3}}\times x^{\frac{1}{3}}=x^{\frac{1}{3}+\frac{1}{3}+\frac{1}{3}}=x^1=x\]

OpenStudy (anonymous):

if you like exponential notation, then \[\sqrt[3]{x^3}=x^{\frac{3}{3}}=x^1=x\]

OpenStudy (anonymous):

\[\frac{1}{x^{-1}}\] the minus one in the exponent means take the reciprocal (flip it) and the reciprocal of \(\frac{1}{x}\) is \(x\)

OpenStudy (anonymous):

did u just see what i wrote a couple minutes ago?

OpenStudy (anonymous):

I know they all equal x I've already solved for all of them but Idk how to highlight the names of the properties of exponents.

OpenStudy (anonymous):

ok for the one that says \[\large x^{\frac{1}{3}}\times x^{\frac{1}{3}}\times x^{\frac{1}{3}}=x^{\frac{1}{3}+\frac{1}{3}+\frac{1}{3}}=x^1=x\] the property is that when you multiply you add the exponents

OpenStudy (anonymous):

for the one that says \[\frac{1}{x^{-1}}\] the property is that \[b^{-1}=\frac{1}{b}\] i.e. the reciprocal, and the reciprocal of \(\frac{1}{x}\) is \(x\)

OpenStudy (anonymous):

\[\sqrt[11]{x^5x^4x^2}=\sqrt[11]{x^{11}}=x\] the property is 1) when you multiply you add the exponents and 2) \(\sqrt[n]{x^n}=x\)

OpenStudy (anonymous):

same property that explains why \[\sqrt[3]{x^3}=x\]

OpenStudy (anonymous):

tell your friend to stop emailing with such stupid questions

OpenStudy (anonymous):

LOOOOOL!! ikr! who e-mails stuff like that these days anyways!

OpenStudy (anonymous):

no one

OpenStudy (anonymous):

hahaha seriously!

OpenStudy (anonymous):

is it okay if you help me explain how I got all the same answer in complete sentences explaining properties as well plz?!

OpenStudy (anonymous):

i thought i did above do you see the explanation in each answer?

OpenStudy (anonymous):

to recap 1) \(\sqrt[n]{x^n}=x\) for the the first one in english, the nth root of x to the nth power is x

OpenStudy (anonymous):

yeah but in complete sentences

OpenStudy (anonymous):

2) when you multiply with the same base, you add the exponents in the case one third plus one third plus one third is one and anything to the power of one is itself

OpenStudy (anonymous):

its one third TIMES

OpenStudy (anonymous):

3) the property is that raising a number or variable to the power of minus one means take the reciprocal the reciprocal of one over x is x

OpenStudy (anonymous):

not plus lol

OpenStudy (anonymous):

yes, what i said when you MULTIPLY you ADD the exponents

OpenStudy (anonymous):

OH okayyy!!

OpenStudy (anonymous):

and the last one \[\sqrt[11]{x^5x^4x^2}=\sqrt[11]{x^{11}}=x\] is a combination of two properties when you MULTIPLY you ADD the exponents, and the nth root of x to the nth power is x

OpenStudy (anonymous):

can u write however ur explaining right now for every single one of them plz?

OpenStudy (anonymous):

i meant "i wrote them all in english, one at a time"

OpenStudy (anonymous):

also labelled them 1), 2), 3), 4)

OpenStudy (anonymous):

Thank you!!!!!!!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ur awesome!